OpenStudy (anonymous):

If C = 5/9(F-32) is the formula for Celsius, What is the formula for Fahrenheit?

OpenStudy (anonymous):

You should make F the subject of the formula

OpenStudy (anonymous):

So should I do this F = 5/9(C) + 32

OpenStudy (anonymous):

Yes But for more information you can search in google

OpenStudy (anonymous):

Never Mind that I got the answer

OpenStudy (anonymous):

It is better to use this method\[\frac{ T^{0}c-0^{0} }{ 100^{0}-0^{0} }=\frac{ T^{0}c -32}{ 212-32 }=\frac{ T^{0}k-273 }{ 373-273 }\]The numbers on the numerator are freeing point in their respective measure and the numbers on the denominator are boiling points, of water. NB Rearranging the ration will give u the usual formulas